# Self-Rewarding Large Vision-Language Models for Optimizing Prompts in Text-to-Image Generation

### v1

@@ Abstract
Hongji Yang      Yucheng Zhou      Wencheng Han      Jianbing Shen11footnotemark: 1 SKL-IOTSC, CIS, University of Macau yc47942@um.edu.mo, yucheng.zhou@connect.um.edu.mo wencheng256@gmail.com, jianbingshen@um.edu.mo
@@ 1 Introduction
Text-to-image models Rombach et al. (2022); Saharia et al. (2022); Yu et al. (2022) can generate diverse high-quality images based on user-provided prompts. However, effective interaction with these models requires users to possess specific expertise, including familiarity with specialized vocabularies, e.g., “35mm” for camera parameters and “art by Greg Rutkowski” to invoke a particular artistic style. As shown in Figure 1, rewriting prompts according to model-specific knowledge significantly improves the quality of the generated images.
@@ 2.2 Self-Rewarding Learning
[…]
@@
Assuming a pair of responses $<\textbf{y}^{1},\textbf{y}^{2}>$, with a human annotator labeling one of them to be more aligned with human preferences, denoted as $\textbf{y}^{w}\succ\textbf{y}^{l}|x$. the Bradley-Terry (BT) Brown et al. (2020) model stipulates that the human preference distribution $p^{*}$ can be written as follows:
@@
Related Work.
The related work can be found in Appendix A.
@@ 3.1 Dataset Construction Pipeline
Prompt rewrite data.
To equip the LVLM with initial rewriting capabilities, rewrite prompts data is first given to train in the supervised fine-tuning (SFT) manner. This type of data is presented in pairs, i.e., {raw prompt, rewritten prompt}, to guide the LVLM to rewrite prompt. Prompt rewrite data contains 104,487 prompt pairs in Cao et al. (2023).
Evaluation data.
The LVLM specializes in image understanding and analyzing Zhu et al. (2024); Li et al. (2022). However, image aesthetics is still a subjective understanding and not easy for models. Although unnecessary, LVLM appears more difficult to score images and construct preference pairs without fine-tuning on evaluation data Yuan et al. (2024b). Therefore, evaluation data is constructed to offer the model some examples to understand the aesthetics. The prompt data in the previous paragraph is used to generate images by diffusion model Rombach et al. (2022). These images are then scored using the prompt template in the LVLM-as-a-judge way Zheng et al. (2024); Chen et al. (2024a). We use a subset of the aesthetics training set and the Pickapic-v1 training set. The paired images and evaluation prompt are both fed into VILA-40B as well as the facts to obtain the response as truth response for training. We add a truth cue to the evaluation prompt, if the VILA-40B’s scoring of the two images is consistent with the facts, the evaluation response is retained.
Judgment data.
To improve the LVLM’s evaluation capabilities for responses Zheng et al. (2024); Zhu et al. (2023), we also create a dataset for the model to understand which responses to aesthetics or alignment are more reasonable. Specifically, we construct data that assign higher scores to reasonable responses and lower scores to irrelevant responses, which is some kind of confidence. To improve the evaluation ability of LVLM, we create a new dataset to evaluate the model’s responses to aesthetics and PickScore of the images. Similarly, we feed the standard (i.e., VILA-40B’s response) and irrelevant answers into VILA-40B, and if the model has a higher score in the standard answer than in the irrelevant answer, the answer remains in the dataset. A common observation is that most images generated from rewritten prompts are typically assigned scores of 3 or 4. Therefore, we also include some raw prompts to maintain a more uniform score distribution.
@@ 3.3.2 Training on Preference Assessing
where $\theta_{t}$ denotes the $t$-th parameters of the model, the $r(\textbf{x},\textbf{y})$ and the $E(\textbf{y},\textbf{R})$ denote the reward function of the prompt $x$ and the response $y$, respectively. Since the self-reward model is an iterative model, in which the iterative process results in a series of models with different weights of the same structure. For better understanding, the parameter of the result model by Equation 5 is denoted as $\theta_{0}$, while the parameters of the result models optimized through Equation 8 are denoted as $\{\theta_{1},\theta_{2},\cdots\}$. As the model’s parameters continue to be optimized for the ability to follow instructions, so does the reward ability, thereby facilitating self-improvement.
@@ 3.4 RL from AI Feedback
AI Feedback.
One of the main investigations of our work is how to use prior knowledge of large models to guide text-to-image models. Unlike time-consuming and labour-intensive human feedback training, AI feedback Lee et al. (2023) training can provide reward signals for a given task through its own knowledge. This approach usually requires two models, one acting as a solver of the downstream task, and one acting as a judge of the solver. It is feasible to apply an external language model as a judge (or reward model) or to use only itself (i.e. a model that acts as both a solver and a judge) to achieve self-improvement in a specific task.
Therefore, preference pairs $\langle$ raw prompt $x$, winner prompt $y^{w}$, loser prompt $y^{l}$rangle need to be constructed to train the model. The model first generates different $N$ candidates from randomly selected raw prompts (in the previous data). After generating images using a fixed text-to-image model, the images and evaluation prompts are input into the model for scoring. Next, the prompts corresponding to the highest and lowest-scored images are treated as winners and losers, respectively. In addition, the image evaluation responses are judged by LVLM and the pairwise data for DPO are also constructed. The model is then tuned with DPO Rafailov et al. (2024).
DPO.
Assuming access to a static dataset of comparisons $\mathcal{D}=\{\textbf{x}_{i},\textbf{y}^{w}_{i},\textbf{y}^{l}_{i}\}$, which is sampled from $p^{*}$. The optimal RLHF policy $\pi^{*}$ under the Bradley-Terry model satisfies the preference model:
@@
[…]
@@
3.5 LVLM-as-a-judge for Aesthetics and Re Rewarding
@@
Aesthetic Score.
Aesthetics are assessed by prompting the model to consider aspects such as composition, color, lighting, and visual appeal. To facilitate judgment-making, we design a grading system that allows the model to attribute a certain grade to the generated images, thus obtaining the corresponding score. For each evaluation, the model is asked to provide both the score and a brief explanation, employing a chain-of-thought reasoning approach.
PickScore Kirstain et al. (2024).
PickScore is an important metric to measure human preferences. We consider whether the image represents the given text in a way that is favored by humans. Nonetheless, it is difficult for the model to understand human preferences directly, so we included evaluation data (details in Section 3.2) for model initialization.
Relevance Score.
The judgment template for the relevance score uses the same hierarchical form as the aesthetic score. The key difference is that it evaluates the model’s attention to the user instructions for text-to-image generation in three areas: presence of the object, accurate count of the object, and correct relationships between objects.
@@ 4.1 Settings
Experiments are conducted on the public text-to-image model Stable Diffusion v1.5, and the denoising steps are set to 20 to accelerate the image sampling. The base model we employ in this paper is the smaller model VILA-v1.5-8b Lin et al. (2024). For SFT, we use AdamW optimizer ($\beta_{1}=0.9,\beta_{2}=0.95$), with a batch size of 16 and a weight decay of 0.1. We use an initial learning rate of 2e-5, with a linear warm-up and cosine decay schedule. To improve the diversity of the candidates, we sample $N=8$ candidates with temperature $T=0.9,p=0.9$ from one raw prompt. When validating these candidates with the reward model, we utilize a temperature $T=0.9,p=0.9$, sampling three times and averaging the scores to determine the final score for further DPO training. The overall score is calculated by summing up the aesthetics score, pick Score and the relevance score. The highest one and the lowest one are kept as winners and losers but discarded if they have the same score. We perform two DPO iterations and the size of each prompt rewrite training data is 10k and 20k, respectively. Besides, 10k judgment data is employed to improve image evaluation. In the stage of DPO, we employ AdamW with an initial learning rate of 1e-5, 5e-6 without weight decay. The batch size is set to 32 and $\beta$ in DPO is set to 0.1. The model is trained for 4 epochs at each iteration. All experiments are conducted on 4×4$\times$4 × NVIDIA A800 80G GPUs. For evaluation, we adopt beam search Vijayakumar et al. (2016) with a beam size of 4 and a length penalty of 1.0.
@@
$\operatorname*{LVLM}_{\operatorname*{SFT}}$ 20.79 5.95 0.25
@@
Table 1: Evaluation of the aesthetic score and CLIP score on Beautiful-Prompt test set. “${\mathbb{D}}_{rl}$” means the size of the train set we used in reinforcement learning. “Aes.” and “CLIP” mean the aesthetic and CLIP scores, respectively.
@@
[…]
@@ 4.4 Ablation Study
We perform ablation experiments from two perspectives: how the prompt affects the LVLM as a reward model and how self-rewarding training improves the model. First, as can be seen in Table 3, the model without Pick Prompt for rewarding received a high aesthetic score but leads a significant drop in PickScore. It also suggests that some shifts exist between human preferences and image aesthetics. In addition, it can be observed a reduction of the aesthetic score and CLIP score when missing the aesthetic prompt or relevance prompt. Second, we compare the performance between the model using self-rewarding and fixed-rewarding. Self-rewarding outperforms the fixed-rewarding methods. More discussion about the self-rewarding method can be found in the Appendix E.
@@ 4.7 Quantitative Results
We present more visual results in Figure 5, it can be observed that our method rewrites the main content appropriately, and describes the detail, e.g., art style, light, etc. The images generated by these optimized prompts have better aesthetics than the images generated from the original prompts, which are bland visually. In addition, some incorrect prompts can also be corrected. For example, the prompt “Astronaut rides horse" mislead the text-to-image models to generate an astronaut standing beside a horse. However, our approach rewrites it as “…an astronaut riding a horse…" to make the image more in line with the user’s intention. In addition, the model expands some descriptions on some objects, like “futuristic"" relative to the “in 2456”. More results are in the Appendix D.
@@ 5 Conclusion
Appendix A Related Work
Prompt Engineering.
Hao et al. (2024) propose a prompt adaption framework for prompt engineering. To implement reinforcement learning fine-tuning, a reward function for image aesthetics and alignment is defined. Bestprompt Pavlichenko and Ustalov (2023) is proposed to detect keywords by genetic algorithm and then form prompts to obtain the better aesthetics of images. Beautiful-Prompt Cao et al. (2023) first trains two reward models: Aesthetics and PickScore with a large dataset and then optimizes the language model using PPO. NeuroPrompt Rosenman et al. (2023) utilizes constrained text decoding with a pre-trained language model to produce prompts. Datta et al. (2023) is proposed as a Prompt Expansion framework to improve the diversity in text-to-image generation. Inspired by visual language modeling, recently, more and more methods Li et al. (2022); Liu et al. (2024); Chen et al. (2024a); Yang et al. (2023); Yu et al. (2023) attempt to use a prior knowledge of LVLM to analyze images. The vision-language model is a large multimodal model, which bridges the gap between language and images. CLIP Radford et al. (2021) is trained with large-scale paired text and images using contrastive learning. In this way, the pre-trained LVLMs capture rich vision-language correspondence knowledge. ALIGN Jia et al. (2021) scales up the training process, using the larger images-text pairs but noisy data. Recently, with the great success of Large Language Models (LLMs), some work has been devoted to enabling LLMs to use image inputs. OpenFlamingo Awadalla et al. (2023) and LLaMA-Adapter Zhang et al. (2023) construct multimodal models based on the best LLM LlaMA Touvron et al. (2023). To further improve the model’s instruction-following abilities, LlaVA Liu et al. (2024) employs visual instruction tuning that yields promising results. ViLA Lin et al. (2024) achieves multi-image reasoning through a better training strategy.
Text-to-Image generation.
Text-to-image models usually refer to the generative model which synthesizes an image from a given text. Earlier work, GAN Reed et al. (2016); Tao et al. (2022) and VAE Ramesh et al. (2021); Ding et al. (2021) have been extensively studied in this field. Recently, the diffusion-based models Rombach et al. (2022); Gu et al. (2022) further improve the quality of text-to-image generation. Given the generated output is no longer a serious concern, exploring how to optimize the prompt to maximize the potential of generative models becomes the primary object of our research. Several works Hao et al. (2024); Cao et al. (2023) have been put into unlocking the maximum potential of text-to-image models by optimizing prompts using large language models. Besides, some reinforcement learning-based methods have gained consistent improvements in changing the output preferences of large language models. Reinforcement Learning from Human Feedback (RLHF) Yuan et al. (2024a) yields promising results in various NLP tasks Christiano et al. (2017); Ibarz et al. (2018); Stiennon et al. (2020); Jaques et al. (2019), at the cost of collecting large amounts of human feedback. Proximal Policy Optimization (PPO) Schulman et al. (2017) is optimized according to the reward signal rather than human feedback and remains stable by approximating the old and new strategies. With these challenges in mind, Direct Preference Optimization (DPO) Rafailov et al. (2024) avoids the drawback of large amounts of human-annotated data to train reward models by using its language model as a reward model. The different strategies to provide rewards are shown in Figure 6.
Appendix B Comparisons with the common tags

### v2

@@ Abstract
Hongji Yang, Yucheng Zhou, Wencheng Han, Jianbing Shen [^1] Corresponding author. This work was supported by the Science and Technology Development Fund of Macau SAR (FDCT) under grants 0102/2023/RIA2 and 0154/2022/A3 and 001/2024/SKL and CG2025-IOTSC, the National Natural Science Foundation of China (No. 624B2002), and the Jiangyin Hi-tech Industrial DevelopmentZone under the Taihu Innovation Scheme (EF2025-00003-SKL-IOTSC). SKL-IOTSC, CIS, University of Macau yc47942@um.edu.mo, yucheng.zhou@connect.um.edu.mo, wencheng256@gmail.com
@@ 1 Introduction
Text-to-image models Rombach et al. (2022); Saharia et al. (2022); Yu et al. (2022) can generate diverse high-quality images based on user-provided prompts. However, effective interaction with these models requires users to possess specific expertise, including familiarity with specialized vocabularies, e.g., “35mm” for camera parameters and “art by Greg Rutkowski” to invoke a particular artistic style. As shown in Figure˜1, rewriting prompts according to model-specific knowledge significantly improves the quality of the generated images.
@@ 2.2 Self-Rewarding Learning
Self-Rewarding Learning uses the same model to perform iterative training to realize self-improvement. Given a sequence pair $(x,y)$, reinforcement learning fine-tuning demands the definition of the reward function $\textbf{r}(\textbf{x},\textbf{y})$ to quantify the value of the response y to the given input x. The objective of the RL fine-tuning can be defined as:
@@
Assuming a pair of responses $\textbf{y}^{1},\textbf{y}^{2}>$, with a human annotator labeling one of them to be more aligned with human preferences, denoted as $\textbf{y}^{w}\succ\textbf{y}^{l}|x$. the Bradley-Terry (BT) Brown et al. (2020) model stipulates that the human preference distribution $p^{*}$ can be written as follows:
@@
2.3 Related Work
Prompt Engineering. Hao et al. (2024) propose a prompt adaption framework for prompt engineering. To implement reinforcement learning fine-tuning, a reward function for image aesthetics and alignment is defined. Bestprompt Pavlichenko and Ustalov (2023) is proposed to detect keywords by genetic algorithm and then form prompts to obtain the better aesthetics of images. Beautiful-Prompt Cao et al. (2023) first trains two reward models: Aesthetics and PickScore with a large dataset and then optimizes the language model using PPO. NeuroPrompt Rosenman et al. (2023) utilizes constrained text decoding with a pre-trained language model to produce prompts. Datta et al. (2023) is proposed as a Prompt Expansion framework to improve the diversity in text-to-image generation. Inspired by visual language modeling, recently, more and more methods Li et al. (2022); Liu et al. (2024); Chen et al. (2024a); Yang et al. (2023); Yu et al. (2023) attempt to use a prior knowledge of LVLM to analyze images. The vision-language model is a large multimodal model, which bridges the gap between language and images. CLIP Radford et al. (2021) is trained with large-scale paired text and images using contrastive learning. In this way, the pre-trained LVLMs capture rich vision-language correspondence knowledge. ALIGN Jia et al. (2021) scales up the training process, using the larger images-text pairs but noisy data. Recently, with the great success of Large Language Models (LLMs), some work has been devoted to enabling LLMs to use image inputs. OpenFlamingo Awadalla et al. (2023) and LLaMA-Adapter Zhang et al. (2023) construct multimodal models based on the best LLM LlaMA Touvron et al. (2023). To further improve the model’s instruction-following abilities, LlaVA Liu et al. (2024) employs visual instruction tuning that yields promising results. ViLA Lin et al. (2024) achieves multi-image reasoning through a better training strategy.
Text-to-Image generation. Text-to-image models usually refer to the generative model which synthesizes an image from a given text. Earlier work, GAN Reed et al. (2016); Tao et al. (2022) and VAE Ramesh et al. (2021); Ding et al. (2021) have been extensively studied in this field. Recently, the diffusion-based models Rombach et al. (2022); Gu et al. (2022) further improve the quality of text-to-image generation. Given the generated output is no longer a serious concern, exploring how to optimize the prompt to maximize the potential of generative models becomes the primary object of our research. Several works Hao et al. (2024); Cao et al. (2023) have been put into unlocking the maximum potential of text-to-image models by optimizing prompts using large language models. Besides, some reinforcement learning-based methods have gained consistent improvements in changing the output preferences of large language models. Reinforcement Learning from Human Feedback (RLHF) Yuan et al. (2024a) yields promising results in various NLP tasks Christiano et al. (2017); Ibarz et al. (2018); Stiennon et al. (2020); Jaques et al. (2019), at the cost of collecting large amounts of human feedback. Proximal Policy Optimization (PPO) Schulman et al. (2017) is optimized according to the reward signal rather than human feedback and remains stable by approximating the old and new strategies. With these challenges in mind, Direct Preference Optimization (DPO) Rafailov et al. (2024) avoids the drawback of large amounts of human-annotated data to train reward models by using its language model as a reward model. The different strategies to provide rewards are shown in Figure˜2.
@@ 3.1 Dataset Construction Pipeline
Prompt rewrite data. To equip the LVLM with initial rewriting capabilities, rewrite prompts data is first given to train in the supervised fine-tuning (SFT) manner. This type of data is presented in pairs, i.e., {raw prompt, rewritten prompt}, to guide the LVLM to rewrite prompt. Prompt rewrite data contains 104,487 prompt pairs in Cao et al. (2023).
Evaluation data. The LVLM specializes in image understanding and analyzing Zhu et al. (2024); Li et al. (2022). However, image aesthetics is still a subjective understanding and not easy for models. Although unnecessary, LVLM appears more difficult to score images and construct preference pairs without fine-tuning on evaluation data Yuan et al. (2024b). Therefore, evaluation data is constructed to offer the model some examples to understand the aesthetics. The prompt data in the previous paragraph is used to generate images by diffusion model Rombach et al. (2022). These images are then scored using the prompt template in the LVLM-as-a-judge way Zheng et al. (2024); Chen et al. (2024a). We use a subset of the aesthetics training set and the Pickapic-v1 training set. The paired images and evaluation prompt are both fed into VILA-40B as well as the facts to obtain the response as truth response for training. We add a truth cue to the evaluation prompt, if the VILA-40B’s scoring of the two images is consistent with the facts, the evaluation response is retained.
Judgment data. To improve the LVLM’s evaluation capabilities for responses Zheng et al. (2024); Zhu et al. (2023), we also create a dataset for the model to understand which responses to aesthetics or alignment are more reasonable. Specifically, we construct data that assign higher scores to reasonable responses and lower scores to irrelevant responses, which is some kind of confidence. To improve the evaluation ability of LVLM, we create a new dataset to evaluate the model’s responses to aesthetics and PickScore of the images. Similarly, we feed the standard (i.e., VILA-40B’s response) and irrelevant answers into VILA-40B, and if the model has a higher score in the standard answer than in the irrelevant answer, the answer remains in the dataset. A common observation is that most images generated from rewritten prompts are typically assigned scores of 3 or 4. Therefore, we also include some raw prompts to maintain a more uniform score distribution.
@@ 3.3.2 Training on Preference Assessing
where $\theta_{t}$ denotes the $t$-th parameters of the model, the $r(\textbf{x},\textbf{y})$ and the $E(\textbf{y},\textbf{R})$ denote the reward function of the prompt $x$ and the response $y$, respectively. Since the self-reward model is an iterative model, in which the iterative process results in a series of models with different weights of the same structure. For better understanding, the parameter of the result model by Equation˜5 is denoted as $\theta_{0}$, while the parameters of the result models optimized through Equation˜8 are denoted as $\{\theta_{1},\theta_{2},\cdots\}$. As the model’s parameters continue to be optimized for the ability to follow instructions, so does the reward ability, thereby facilitating self-improvement.
@@ 3.4 RL from AI Feedback
AI Feedback. One of the main investigations of our work is how to use prior knowledge of large models to guide text-to-image models. Unlike time-consuming and labour-intensive human feedback training, AI feedback Lee et al. (2023) training can provide reward signals for a given task through its own knowledge. This approach usually requires two models, one acting as a solver of the downstream task, and one acting as a judge of the solver. It is feasible to apply an external language model as a judge (or reward model) or to use only itself (i.e. a model that acts as both a solver and a judge) to achieve self-improvement in a specific task.
Therefore, preference pairs $\langle$ raw prompt $x$, winner prompt $y^{w}$, loser prompt $y^{l}$ $\rangle$ need to be constructed to train the model. The model first generates different $N$ candidates from randomly selected raw prompts (in the previous data). After generating images using a fixed text-to-image model, the images and evaluation prompts are input into the model for scoring. Next, the prompts corresponding to the highest and lowest-scored images are treated as winners and losers, respectively. In addition, the image evaluation responses are judged by LVLM and the pairwise data for DPO are also constructed. The model is then tuned with DPO Rafailov et al. (2024).
DPO. Assuming access to a static dataset of comparisons $\mathcal{D}=\{\textbf{x}_{i},\textbf{y}^{w}_{i},\textbf{y}^{l}_{i}\}$, which is sampled from $p^{*}$. The optimal RLHF policy $\pi^{*}$ under the Bradley-Terry model satisfies the preference model:
@@
Given a preference pair $\langle\textbf{x}$, $\textbf{y}^{w}$, $\textbf{y}^{l}\rangle$, the object function of DPO is to seek a maximum likelihood of the parameterized policy $\pi_{\theta}$ by reference model $\pi_{ref}$.
@@
3.5 LVLM-as-a-judge for Aesthetics and Rewarding
@@
Aesthetic Score. Aesthetics are assessed by prompting the model to consider aspects such as composition, color, lighting, and visual appeal. To facilitate judgment-making, we design a grading system that allows the model to attribute a certain grade to the generated images, thus obtaining the corresponding score. For each evaluation, the model is asked to provide both the score and a brief explanation, employing a chain-of-thought reasoning approach.
PickScore Kirstain et al. (2024). PickScore is an important metric to measure human preferences. We consider whether the image represents the given text in a way that is favored by humans. Nonetheless, it is difficult for the model to understand human preferences directly, so we included evaluation data (details in Section 3.2) for model initialization.
Relevance Score. The judgment template for the relevance score uses the same hierarchical form as the aesthetic score. The key difference is that it evaluates the model’s attention to the user instructions for text-to-image generation in three areas: presence of the object, accurate count of the object, and correct relationships between objects.
@@ 4.1 Settings
Experiments are conducted on the public text-to-image model Stable Diffusion v1.5, and the denoising steps are set to 20 to accelerate the image sampling. The base model we employ in this paper is the smaller model VILA-v1.5-8b Lin et al. (2024). For SFT, we use AdamW optimizer ($\beta_{1}=0.9,\beta_{2}=0.95$), with a batch size of 16 and a weight decay of 0.1. We use an initial learning rate of 2e-5, with a linear warm-up and cosine decay schedule. To improve the diversity of the candidates, we sample $N=8$ candidates with temperature $T=0.9,p=0.9$ from one raw prompt. When validating these candidates with the reward model, we utilize a temperature $T=0.9,p=0.9$, sampling three times and averaging the scores to determine the final score for further DPO training. The overall score is calculated by summing up the aesthetics score, pick Score and the relevance score. The highest one and the lowest one are kept as winners and losers but discarded if they have the same score. We perform two DPO iterations and the size of each prompt rewrite training data is 10k and 20k, respectively. Besides, 10k judgment data is employed to improve image evaluation. In the stage of DPO, we employ AdamW with an initial learning rate of 1e-5, 5e-6 without weight decay. The batch size is set to 32 and $\beta$ in DPO is set to 0.1. The model is trained for 4 epochs at each iteration. All experiments are conducted on $4\times$ NVIDIA A800 80G GPUs. For evaluation, we adopt beam search Vijayakumar et al. (2016) with a beam size of 4 and a length penalty of 1.0.
@@
$\operatorname*{LVLM}_{\operatorname*{SFT}}$ - 20.79 5.95 0.25
@@
Table 1: Evaluation of the aesthetic score and CLIP score on Beautiful-Prompt test set. ${\mathbb{D}}_{rl}$” means the size of the train set we used in reinforcement learning. “Aes.” and “CLIP” mean the aesthetic and CLIP scores, respectively.
@@
Table 2: Evaluation of the aesthetic score and CLIP score on DiffusionDB. Values marked with $\dagger$} from Hao et al. (2024).
4.2 Comparative Methods
4.3 Results
@@ 4.4 Ablation Study
We perform ablation experiments from two perspectives: how the prompt affects the LVLM as a reward model and how self-rewarding training improves the model. First, as can be seen in Table 3, the model without Pick Prompt for rewarding received a high aesthetic score but leads a significant drop in PickScore. It also suggests that some shifts exist between human preferences and image aesthetics. In addition, it can be observed a reduction of the aesthetic score and CLIP score when missing the aesthetic prompt or relevance prompt. Second, we compare the performance between the model using self-rewarding and fixed-rewarding. Self-rewarding outperforms the fixed-rewarding methods. More discussion about the self-rewarding method can be found in the Appendix D.
@@ 4.7 Quantitative Results
We present more visual results in Figure 6, it can be observed that our method rewrites the main content appropriately, and describes the detail, e.g., art style, light, etc. The images generated by these optimized prompts have better aesthetics than the images generated from the original prompts, which are bland visually. In addition, some incorrect prompts can also be corrected. For example, the prompt “Astronaut rides horse" mislead the text-to-image models to generate an astronaut standing beside a horse. However, our approach rewrites it as “…an astronaut riding a horse…" to make the image more in line with the user’s intention. In addition, the model expands some descriptions on some objects, like “futuristic"" relative to the “in 2456”. More results are in the Appendix C.
@@ 5 Conclusion
Appendix A Comparisons with the common tags